# Downsizing Diffusion Models for Cardinality Estimation

### v1

@@ 2.2. Score-Based Diffusion Models
where $k(t)=e^{\int_{0}^{t}\alpha(s)ds}$, $\sigma^{2}(t)$ satisfies $\frac{d\sigma^{2}}{dt}=\beta(t)-2\alpha(t)\sigma^{2}(t)$ with initial condition $\sigma^{2}(0)=0$, $\tau$ in $\mathbb{R}^{d}$ is the scaling operator $(\tau_{k}\circ f)(x)=k^{d}f(kx)$, and $\varphi_{\sigma^{2}(t)}$ in $\mathbb{R}^{d}$ is the probability density function of the normal distribution $\mathcal{N}(0,\sigma^{2}(t)I_{d})$.
The Backward Process, in turn, generates a sample point by first sampling a starting point $\tilde{x}(T)\sim\mathcal{N}(0,\sigma^{2}(T)I_{d})\approx p_{T,\alpha,\beta}(x)$, then use it as an initial value to solve the reverse diffusion SDE (Song et al., 2021b)
@@
(where $\hat{w}$ denotes a standard Wiener process in the reverse time direction), or its associated probability flow ordinary differential equation (ODE)
$\frac{dx}{dt}=-[\alpha(t)x+\frac{1}{2}\beta^{2}(t)\nabla\log p_{t}(x)].$
Here, $\nabla\log p_{t}(x)$, the only unknown term in both equations, is called the score function. To approximate it, diffusion models train a neural network $s_{\theta}(x,t)$ to minimize the score prediction error
(3) $\int_{0}^{T}\mathrm{E}_{p_{t,appr}(x)}[\beta^{2}(t)\left\|\nabla\log p_{t,appr}(x)-s_{\theta}(x,t)\right\|_{2}^{2}]dt,$
@@
being our ”best guess” for $p_{t}(x)$ from $N$ i.i.d. samples.
The accuracy of score-based diffusion models has been verified both experimentally and theoretically. In fact, the KL divergence between $p_{0}(x)$, the original underlying distribution, and $p_{0,\theta(x)}$, which $\tilde{x}(0)$ is distributed accordingly, would satisfy (Song et al., 2021a)
@@
where the equal sign holds whenever $s_{\theta}(x,t)$ is the exact score function of a diffusion process (not necessarily $\nabla\log p_{t,\alpha,\beta}(x)$).
Noting that the above formula did not require how ”similar” $p_{0}$ must be from $p_{0,\theta}$, replacing $p_{0}(x)$ with $\mathcal{N}(x_{0},\epsilon^{2}I_{d})$, a normal distribution centered at an arbitrary point $x_{0}$, and letting $\epsilon\to 0$ produces a more interesting formula (Song et al., 2021a)
$\log p_{0,\theta}(x_{0})\approx\mathrm{E}_{p_{0T}(x|x_{0})}\log\varphi_{\sigma^{2}(T)}(x)+\int_{0}^{T}n\alpha(t)dt+\frac{1}{2}\int_{0}^{T}\beta^{2}(t)$
$\mathrm{E}_{p_{0t(x|x_{0})}}[\left\|\nabla\log p_{0t}(x|x_{0})-s_{\theta}(x,t)\right\|_{2}^{2}-\left\|\nabla\log p_{0t}(x|x_{0})\right\|_{2}^{2}]dt,$
where $p_{0t}(x|x_{0})=\mathcal{N}(\frac{x_{0}}{k(t)},\sigma^{2}(T))$ is the distribution derived by perturbing $\delta_{x_{0}}$ for time $t$ under the same perturbation scheme. Again, the equal sign would exactly hold when $s_{\theta}(x,t)$ is the score function of some diffusion process, a condition that would almost be met if $s_{\theta}$ can match $\nabla\log p_{t,\alpha,\beta}(x)$ sufficiently well.
The above formula lays the cornerstone for our model, allowing us to estimate a probability density function at any point, and hence conduct cardinality estimation, by training $s_{\theta}(x,t)$ and integrating it across time and sample space. In section 3 and section 4, we shall design the score estimator and density estimator that can finish the task in a matter of milliseconds.
@@ 3.1. Model Choice
The score estimator trains a neural network, $s_{\theta}(x,t)$, to approximate the score function $\{\nabla\log p_{t}(x)|t\in(0,T)\}$, thus providing the density estimator with the data it needs to calculate pointwise densities. Compared to training a network that directly outputs $\nabla\log p_{t}(x)$, works on score-based diffusion models usually find one of the following two methods much more efficient.
These approaches, named the data prediction model (Lu et al., 2025)(Kingma et al., 2021) and the noise prediction model (Song et al., 2021b)(Lu et al., 2025) respectively, are both inspired by a corollary of formula (1) that $x(t)\sim p_{t}(x)$ can be expressed as the sum of two independent variables $x_{sig}$ and $x_{noise}$, where $k(t)x_{sig}\sim p_{0}(x)$ and $\sigma(t)^{-1}x_{noise}\sim\mathcal{N}(0,I_{d})$. Therefore, $\nabla\log p_{t}(x)$ would equal
@@
Conventional works on diffusion models usually stick to one of the two designs throughout the score matching process. However, we find that identifying the best-fit scenario of each model can prove remarkably easy. Indeed, for any choice of $(x,t)$, absolute error from approximating $\mathrm{E}_{p_{s}(x_{sig}|x)}x_{sig}$ and $\mathrm{E}_{p_{n}(x_{sig}|x)}x_{noise}$ equally affect the error term $\left\|\nabla\log p_{t,approx}(x)-s_{\theta}(x,t)\right\|_{2}^{2}$, as $x_{sig}$ and $x_{noise}$ determine each other by a simple plus-and-minus relation. Thus, predicting the smaller of $x_{sig}$ and $x_{noise}$, that is, $x_{noise}$ when $t$ is small and $x_{sig}$ when $t$ is large, naturally leads to a larger relative error tolerance.
With this in mind, we have ADC’s score estimator break score matching into two parts: approximating $\nabla\log p_{t}(x)$ with a data prediction model $s_{\theta_{2},tail}(x,t)$ when $t$ is above a threshold $T_{trunc}$, and a noise prediction model $s_{\theta_{1},head}(x,t)$ when $t$ is below it. To the best of our knowledge, this is the first time such a practice has been adpoted in the field of score-based diffusion models.
Many studies (Song et al., 2023) (Chen et al., 2023) (Oko et al., 2023) show that $p_{t,appr}(x)$, when constructed from i.i.d. samples, grows drastically larger and more volatile as $t\to 0$ and eventually explodes to infinity. As a result, $s_{\theta_{1},head}$ must be somewhat heavy to produce moderately accurate estimations, but $s_{\theta_{2},tail}$ can be very lightweight yet just as precise, especially since when $p_{t}$ gets sufficiently close to $\mathcal{N}(0,\sigma^{2}(t))$, the data prediction model’s added term,
@@
has already completed the bulk of score matching pretty decently. Therefore, our approach of implementing two separate models not only improves precision, but can also speed up computation by calculating the score function with a heavy network only where it is really needed.
Finally, we would like to point out that the extra storage space cost incurred by training two separate models is quite minimal: In all our test scenarios, the extra network used by the data prediction model, $s_{\theta_{2},tail}$, only takes up slightly more than a third of $_{\theta_{1},head}’s s$torage space.
3.2. Model Architecture Improvement: Introducing QuadNet
@@
    with $\delta_{\sigma^{2}}(x_{0},t)$ a residual term of order $o(\frac{1}{\sigma(t)^{1+\epsilon}})$ for all $\epsilon>0$.
The above findings inspire us to present a network structure called Quadnet. That is, for a $d$-dimensional dataset, we have our network $s_{\theta_{1},head}$ generate a vector $(v_{\sigma^{2}},v_{\sigma},v_{1})$ of length $3d$, and output
@@
To illustrate the advantage of Quadnet, we define a toy dataset as of Figure 2, and have a Quadnet of shape $[3,10,25,25,10,6]$ compete against three compare group networks of shape $[3,10,25,25,10,6,2]$ in performing the score matching task for small $t$. Compared to its competitors that possess only the quadratic ($\frac{1}{\sigma^{2}(t)}$), linear ($\frac{1}{\sigma(t)}$, i.e. a traditional noise prediction model), or constant (i.e. directly approximates $\nabla\log p_{t}(x)$) scaling module, Quadnet converges faster, settles to a lower error, and becomes less prone to falling into saddle points than all of its competitors.
@@
3.3. Mitigating Score Blowup
Despite the enhancements above, it is widely observed that the score function would still grow to a point beyond the approximation ability of any neural networks when $t$ becomes too close to zero. Score-based diffusion models used for image generation mitigate this problem by implementing early stopping (Song et al., 2023) (Oko et al., 2023): that is, training the score matching model (i.e., calculating the score approximation loss) and running the reverse diffusion SDE solver on the interval $(\epsilon,T]$ rather than $(0,T]$, and treating $x(\epsilon)$ rather than $x(0)$ as the final sample.
The score estimator in ADC is trained using the same early stopping approach, with $\epsilon$, the early stopping time, selected among values $\{\frac{1}{640},\frac{1}{320},\frac{1}{160}\}$ using a trial-and-error approach (starting from $\frac{1}{640}$ and increasing $\epsilon$ if the MSE loss proves too big). We notice that the best choice for $\epsilon$ increases with dimensionality and volume of the data distribution’s typical set, but more work is still needed to determine the best choice for $\epsilon$ ahead of model construction.
Another natural consequence of score blowup, theoretically coined in the data prediction model case by (Chen et al., 2023), is that the score approximation error when $t$ is small always accounts for the bulk of the entire error term. In fact, in our model where $T$ is set to 3 and $T_{trunc}$ is set to $1/8$,
$\int_{\epsilon}^{\frac{1}{32}}\mathrm{E}_{p_{t}(x)}\left\|\nabla\log p_{t}(x)-s_{\theta}(x,t)\right\|_{2}^{2}dt$
often contributes to more than 60% of (3)’s size. Therefore, a score estimator’s quality can almost entirely be judged by whether it can perform score matching with enough accuracy on the lower end of the evaluation interval.
With this in mind, ADC speeds up score matching by iterating between exploration epoches and adjustment epoches in its training process. In exploration epoches, the model is told to solely focus on reducing approximation error in the small-$t$ interval ($\epsilon,\frac{1}{32}]$ for $s_{\theta_{1},head}$ and $T_{trunc},\frac{1}{2}]$ for $s_{\theta_{2},tail}$), ignoring the rest of the training set. In adjustment approaches, the model, having learned the most difficult, complicated features of the score function, is now allowed to see the full picture, minimizing (3) in its unedited, unsliced form for most optimal performance. On all experiment datasets in section 6, this approach is always able to speed up training, and sometimes results in a lower final error by escaping a local minimum that can trap a model trained with only adjustment epoches.
@@ 4.1. Model Design
Relying on score information, $s_{\theta}(x,t)$, as input, the density estimator calculates pointwise joint densities for the selectivity estimator using the formula
$\log p_{0,\theta}(x_{0})\approx\mathrm{E}_{p_{0T}(x|x_{0})}\log\varphi_{\sigma^{2}(T)}(x)+\int_{0}^{T}n\alpha(t)dt+\frac{1}{2}\int_{0}^{T}\beta^{2}(t)$
$\mathrm{E}_{p_{0t(x|x_{0})}}[\left\|\nabla\log p_{0t}(x|x_{0})-s_{\theta}(x,t)\right\|_{2}^{2}-\left\|\nabla\log p_{0t}(x|x_{0})\right\|_{2}^{2}]dt,$
where we denote
@@
as $g_{x_{0}}(x,t,s_{\theta})$ for the sake of convenience. Here,
@@
can be exactly calculated with an analytic expression, so our main job would be to estimate
@@
using a numerical integration scheme.
@@
more solidly backed by SDE theory.
ADC chooses the third approach despite the added perturbation, both because it’s easier to analyze theoretically, and because the first two approaches seem to introduce arbitrary spikes in the approximated joint distribution function that significantly reduces the accuracy of the current selectivity estimator, designed using a predictor-corrector importance sampling algorithm. Whether the first two approaches might work for a different selectivity estimator design can be a topic of future research.
@@
To account for such, we present a hybrid integration approach: for integration across time, ADC implements an adaptive stepsize Midpoint Rule Integrator, which is more suited for low-dimensional functions and can be easily modified to capitalize our knowledge of the score function’s behavior; for calculating the expectation across space, ADC implements a Quasi Monte Carlo (Leobacher and Pillichshammer, 2014) integrator much more robust against the curse of dimensionality.
@@
where $0=t_{0}<t_{1}<...<t_{N-1}<t_{N}=T$ is a partition of $[0,T]$, $\tilde{t}_{i}=\frac{t_{i-1}+t_{i}}{2}$; $\{z_{j}\}_{j=1}^{2^{k}}$ is a Sobol sequence of length $2^{k}$ in the unit cube; $\{y(\tilde{t}_{i})\}_{i=1}^{N}$ are $N$ random samples from $\mathrm{U}([0,1]^{d})$, used to perturb $\{z_{j}\}$ so that the integrator will not focus too much on any specific region; $\Phi$ is the distribution function of the standard normal distribution, acting pointwise on each coordinate; and the sum between $z_{j}$ and $y(\tilde{t}_{i})$ is taken in $\mathrm{T}^{n}$ rather than $\mathbb{R}^{n}$ to prevent out-of-domain inputs for $\Phi^{-1}$.
@@ 4.3. Parameter Choice
holds for all $\{x_{i}\}_{i=1}^{M}$ in the database. Meanwhile, one can easily deduce from (5) that
@@
    The partition is dense enough that for the same $y$, fluctuations of $\Delta t_{i}\delta_{t,space}(x_{0},y)$ for different $t$ in $t_{i-1},t_{i}]$ is negligible compared to the total integration error
@@
    The error terms $\{\delta_{i,time}(x_{0})\}_{i=1}^{N}$ are negligible compared to $\{\delta_{\tilde{t}_{i},space}(x_{0},y(\tilde{t_{i}}))\}_{i=1}^{N}$
@@
The decay rate of the second term is estimated to be no slower than $\mathcal{O}(\frac{1}{\sigma^{8}(t)k^{4}(t)})$ when $t\to\infty$, as we already discussed. For the first term, Theorem 3.1 proves that for all $x_{0}$ such that $p_{0}(x_{0})\neq 0$, $\mathrm{E}_{p_{0t}(x|x_{0})}\left\|\nabla\log p_{t+\epsilon}(x)\right\|_{2}^{4}$ can be upper bounded by
@@
which, using the bounds on $\delta_{\sigma^{2}}(x,t)$, can be roughly estimated as proportionate to $\frac{1}{\sigma^{4}(t+\epsilon)}$. Therefore, Cauchy’s Inequality proves the growth rate of
@@
to be no faster than $\mathcal{O}(\frac{1}{\sigma^{2}(t)\sigma^{2}(t+\epsilon)})$ as $t\to 0$, thus justifying our pseudocode implementation.

### v2

@@ 2.2.2. Forward and Backward process
where $\tau$ in $\mathbb{R}^{d}$ is the scaling operator $(\tau_{k}\circ f)(x)=k^{d}f(kx)$ and $\varphi_{\sigma^{2}(t)}$ in $\mathbb{R}^{d}$ is the probability density function of $\mathcal{N}(0,\sigma^{2}(t)I_{d})$.
The Backward Process, in turn, generates a sample point by first sampling a starting point
$\tilde{x}(T)\sim\mathcal{N}(0,\sigma^{2}(T)I_{d})\approx p_{T,\alpha,\beta}(x),$
then using it as an initial value to solve the reverse diffusion SDE (refone)
@@
where $\hat{w}$ denotes a standard Wiener process in reverse time.
To approximate the unknown score function, $\nabla\log p_{t}(x)$, diffusion models train a neural network $s_{\theta}(x,t)$ to minimize
(5) $\int_{0}^{T}\mathrm{E}_{p_{t,appr}(x)}[\beta^{2}(t)\left\|\nabla\log p_{t,appr}(x)-s_{\theta}(x,t)\right\|_{2}^{2}]dt$
@@
being our empirical approximation for $p_{t}(x)$ from $N$ i.i.d. samples.
2.2.3. Error Analysis and Likelihood Evaluation
Theorem 2.2, discovered by Song and Durkan(reftwo), provides a robust error analysis for diffusion models and a formula for estimating pointwise density values of the underlying distribution.
Theorem 2.2.
(reftwo) Let $s_{\theta}(x,t):\mathbb{R}^{d}\times\mathbb{R^{+}}\to\mathbb{R}^{d}$ be a neural network and $p_{0}$ be a probability distribution, and $p_{0,\theta}(x)$ to be the marginal distribution of $\tilde{x}(0)$ after $\tilde{x}(T)\sim\mathcal{N}(0,\sigma^{2}(T)I_{d})$ is perturbed by the reverse diffusion SDE
$dx=-[\alpha(t)x+\beta^{2}(t)s_{\theta}(x,t)]dt+\beta(t)d\hat{w}$
from time $t$ to time 0. Using the notation of Theorem Theorem 2.1, the KL divergence between $p_{0}$ and $p_{\theta}$ can be upper bounded by
@@ 2.2.3. Error Analysis and Likelihood Evaluation
while the value of $\log p_{\theta}(x_{0})$ can be lower bounded by
$\log p_{0,\theta}(x_{0})\geq\mathrm{E}_{p_{0T}(x|x_{0})}\log\varphi_{\sigma^{2}(T)}(x)+\int_{0}^{T}n\alpha(t)dt-\frac{1}{2}\int_{0}^{T}\beta^{2}(t)\cdot$
(8) $\mathrm{E}_{p_{0t(x|x_{0})}}$ $[\left\|\nabla\log p_{0t}(x|x_{0})-s_{\theta}(x,t)\right\|_{2}^{2}-\left\|\nabla\log p_{0t}(x|x_{0})\right\|_{2}^{2}]dt$
with the equal sign holding in both inequalities if there exists a probability distribution $q$ such that $s_{\theta}(x,t)=\nabla\log q_{t}(x),\forall t\in[0,T]$.
Theorem 2.2 lays the foundation for our model, allowing us to estimate a probability density function at any point, and hence conduct cardinality estimation, by training $s_{\theta}(x,t)$ and integrating it across time and sample space.
Our model, ADC, is composed of three key modules: the Score Estimator, the Density Estimator, and the Selectivity Estimator. Sections 3 to 5 shall each be devoted to one of the estimators above.
@@ 3.1. Model Choice
The score estimator trains a neural network, $s_{\theta}(x,t)$, to approximate the score function $\{\nabla\log p_{t}(x)|t\in(0,T]\}$, thus providing the density estimator with the data it needs to calculate pointwise densities. Compared to training a network that directly outputs $\nabla\log p_{t}(x)$, works on score-based diffusion models usually find one of the following two methods much more efficient.
These approaches, named the data prediction model (reftwenty)(refthree) and the noise prediction model (refthree)(refone) respectively, are both inspired by a corollary of (2) that $x(t)\sim p_{t}(x)$ is the sum of two independent variables $x_{sig}$ and $x_{noise}$, where $k(t)x_{sig}\sim p_{0}(x)$ and $\sigma(t)^{-1}x_{noise}\sim\mathcal{N}(0,I_{d})$. Therefore, $\nabla\log p_{t}(x)$ equals
@@
Conventional works on diffusion models usually stick to one of the two designs throughout the score matching process. However, we can easily predict which design works better at a given time $t$. Indeed, as $x=x_{sig}+x_{noise}$, absolute error from approximating $\mathrm{E}_{p_{s}(x_{sig}|x)}x_{sig}$ and $\mathrm{E}_{p_{n}(x_{sig}|x)}x_{noise}$ equally affect the error term $\left\|\nabla\log p_{t,approx}(x)-s_{\theta}(x,t)\right\|_{2}^{2}$ for any choice of $(x,t)$. Thus, predicting the smaller of $x_{sig}$ and $x_{noise}$ would reduce the denominator for relative error, thereby leading to a larger relative error tolerance.
A common property of the forward diffusion process is that the signal-to-noise ratio, ${\mathrm{E}(x_{sig})}/{\mathrm{E}(x_{noise})}$, decreases uniformly with $t$. Thus, we have ADC’s score estimator break score matching into two parts: approximating $\nabla\log p_{t}(x)$ with a data prediction model $s_{\theta_{2},tail}(x,t)$ when $t$ is above a threshold $T_{trunc}$, and a noise prediction model $s_{\theta_{1},head}(x,t)$ when $t$ is below it. To the best of our knowledge, this is the first time such a practice has been adopted in the field of score-based diffusion models.
Many studies (reftwentyseven) (refnineteen) (reffour) show that $p_{t,appr}(x)$, when constructed from i.i.d. samples, grows drastically larger and more volatile as $t\to 0$ and eventually explodes to infinity. As a result, $s_{\theta_{1},head}$ must be somewhat heavy to produce moderately accurate estimations, but $s_{\theta_{2},tail}$ can be very lightweight yet just as precise, especially since when $p_{t}$ gets sufficiently close to $\mathcal{N}(0,\sigma^{2}(t))$, the data prediction model’s added term,
@@
has already completed the bulk of score matching pretty decently. Therefore, our approach of implementing two separate models not only improves precision but can also speed up computation by calculating the score function with a heavy network only where it is really needed. Finally, the extra storage space cost incurred by training two separate models is quite minimal: On all benchmark datasets, the extra network used by the data prediction model, $s_{\theta_{2},tail}$, only takes up roughly a third of $_{\theta_{1},head}’s s$torage space.
3.2. Model Architecture Improvement: QuadNet
@@
    with $\delta_{\sigma^{2}}(x_{0},t)$ a term of order $o(\frac{1}{\sigma(t)^{1+\epsilon}})$ for all $\epsilon>0$.
Such estimates inspire us to present Quadnet, a new network structure. That is, for a $d$-dimensional dataset, we have our network $s_{\theta_{1},head}$ generate a vector $(v_{\sigma^{2}},v_{\sigma},v_{1})$ of length $3d$, and output
@@
To illustrate the advantage of Quadnet, we define a toy dataset as of Figure 2, and have a Quadnet of shape $[3,10,25,25,10,6]$ compete against three compare group networks of shape $[3,10,25,25,10,$ $6,2]$ in performing score matching for small $t$. Compared to its competitors that possess only the quadratic (${1}/{\sigma^{2}(t)}$), linear (${1}/{\sigma(t)}$, i.e. a traditional noise prediction model), or constant (i.e. one that directly approximates $\nabla\log p_{t}(x)$) scaling module, Quadnet converges faster, settles to a lower error, and becomes less prone to falling into saddle points than all of its competitors.
@@
3.3. Choosing Early Stopping Time
Despite the enhancements above, it is widely observed that the score function would still grow to a point beyond the approximation ability of any neural networks when $t$ becomes too close to zero. Score-based diffusion models used for image generation mitigate this problem by implementing early stopping (refnineteen)(reffour): that is, training the score matching model (i.e., calculating the score approximation loss) and running the reverse diffusion SDE solver on the interval $(\epsilon,T]$ rather than $(0,T]$, and treating $x(\epsilon)$ rather than $x(0)$ as the final sample.
The score estimator in ADC is trained using the same early stopping approach, with $\epsilon$, the early stopping time, selected among values $\{\frac{1}{1280},\frac{1}{640},\frac{1}{320},\frac{1}{160}\}$ using a trial-and-error approach (starting from $\frac{1}{1280}$ and increasing $\epsilon$ if the MSE loss proves too big). We notice that the best choice for $\epsilon$ increases with dimensionality and volume of the data distribution’s typical set, but more work is still needed to determine the best choice for $\epsilon$ ahead of model construction.
@@ 4.1. Model Design
Relying on $s_{\theta}(x,t)\approx\nabla\log p_{t}(x)$, as input, the density estimator calculates pointwise joint densities using the formula (15), in which we denote
@@
as $g_{x_{0}}(x,t,s_{\theta})$ for convenience. In the formula (15),
@@
can be calculated analytically, thus leaving the estimation of
@@
to be the bulk of our work.
@@
more numerically stable.
ADC chooses the third approach despite the added perturbation, both because it’s easier to analyze theoretically, and because the first two approaches seem to introduce arbitrary spikes in the approximated joint distribution function that significantly reduce the accuracy of the current selectivity estimator, designed using a predictor-corrector importance sampling algorithm. Whether the first two approaches might work for a different selectivity estimator design can be a topic of future research.
@@
To account for such, we present a hybrid integration approach: for integration across time, ADC implements an adaptive stepsize Midpoint Rule Integrator, which is more suited for low-dimensional functions and can be easily modified to capitalize our knowledge of the score function’s behavior; for calculating the expectation across space, ADC implements a Quasi Monte Carlo (reftwentyone) integrator much more robust against the curse of dimensionality.
@@
where $0=t_{0}<t_{1}<...<t_{N-1}<t_{N}=T$ is a partition of $[0,T]$, $\tilde{t}_{i}={(t_{i-1}+t_{i})}/{2}$; $\{z_{j}\}_{j=1}^{2^{k}}$ is a Sobol sequence of length $2^{k}$ in the unit cube; $\{y(\tilde{t}_{i})\}_{i=1}^{N}$ are $N$ random samples from $\mathrm{U}([0,1]^{d})$, used to perturb $\{z_{j}\}$ so that the integrator will not focus too much on any specific region; $\Phi$ is the distribution function of the standard normal distribution, acting pointwise on each coordinate; and the sum between $z_{j}$ and $y(\tilde{t}_{i})$ is taken in $\mathrm{T}^{n}$ rather than $\mathbb{R}^{n}$ to prevent out-of-domain inputs for $\Phi^{-1}$.
@@ 4.3. Parameter Choice
holds for all $\{x_{i}\}_{i=1}^{n}$. Meanwhile, one can deduce from (Theorem 2.2) that
@@
    The partition is dense enough that for the same $y$, fluctuations of $\Delta t_{i}\delta_{t,space}(x_{0},y)$ for different $t$ in $t_{i-1},t_{i}]$ is negligible compared to the total integration error,
@@
    The error terms $\{\delta_{i,time}(x_{0})\}_{i=1}^{N}$ are negligible compared to $\{\delta_{\tilde{t}_{i},space}(x_{0},y(\tilde{t_{i}}))\}_{i=1}^{N}$, and
@@
The decay rate of the second term is estimated to be no slower than ${O}({1}/(\sigma^{8}(t)k^{4}(t)))$ when $t\to\infty$, as we already discussed. For the first term, Theorem 3.1 proves that for all $x_{0}$ such that $p_{0}(x_{0})\neq 0$, $\mathrm{E}_{p_{0t}(x|x_{0})}\left\|\nabla\log p_{t+\epsilon}(x)\right\|_{2}^{4}$ can be upper bounded by
@@
which, using the bounds on $\delta_{\sigma^{2}}(x,t)$, can be roughly estimated as proportionate to $\frac{1}{\sigma^{4}(t+\epsilon)}$. Therefore, Cauchy’s Inequality proves
@@
to grow at a rate no faster than ${O}({1}/({\sigma^{2}(t)\sigma^{2}(t+\epsilon)}))$ as $t\to 0$, thus justifying our pseudocode implementation.